I am so lost on how to solve this problem....


A woman deposits $50,000 in a savings account with 4% continuously compounded interest. How many years must she wait until the balance has doubled?

Answer by stanbon(75887)

Formula A=Pe^(rt)
You want A to equal 2P
So 2P=Pe^rt
100,000= 50,000e^(rt)
2=e^rt
2=e^(0.04t)
Take the natural log of both sides to get:
ln 2 = 0.04t
t= (ln 2)/0.04
t= 17.33 years.
Cheers,
Stan H.